OpenStudy (anonymous):

How would I write this arith.sequence in standard form? 9,12,15,18

OpenStudy (anonymous):

The common difference is 3. (12-9 or 18-15). The first term is 9. So the general term is a+ (n-1)d where a = first term d = common difference. So in your example, the general term is 9 + (n-1)(3) = 9 + 3n - 3 = 3n + 6.

OpenStudy (anonymous):

Oh, how simple. Thank you.

OpenStudy (anonymous):

Welcome. Worth to jot down that general term "formula" for an arithmetic sequence.

OpenStudy (anonymous):

One final note...if you have the general term, say 3n+6, as in our last example, you can easily figure out "what would be the 120th term of that sequence". Since you have the general form for the sequence, just substitute n = 120, and you get 3(120) + 6 = 366.

OpenStudy (anonymous):

@Yondaime

OpenStudy (anonymous):

Wow, that is mighty useful to know. Thank you again!!

OpenStudy (anonymous):

Very welcome.